OpenStudy (anonymous):

solve the equation for solutions in the interval [0,2π) by first solving for the trigonometric function. (cot x - √3) (√2 sin x + 1) = 0

myininaya (myininaya):

remember a*b=0 means either a=0 or b=0 or both

myininaya (myininaya):

\[\cot(x)-\sqrt{3}=0 \text{ or } \sqrt{2}\sin(x)+1=0\]

myininaya (myininaya):

\[\cot(x)=\sqrt{3} \text{ or } \sin(x)=\frac{-1}{\sqrt{2}}=-\frac{\sqrt{2}}{2}\]

myininaya (myininaya):

for the fist equation when do we have positive for cot(x) when we are in the quadrant 1(+,+) and in quadrant 3(-,-) cot(x)=cos(x)/sin(x) when does this happen \[\frac{\cos(x)}{\sin(x)}=\frac{\frac{\sqrt{3}}{2}}{\frac{1}{2}}\] or when does this happen \[\frac{\cos(x)}{\sin(x)}=\frac{-\frac{\sqrt{3}}{2}}{-\frac{1}{2}}\]

myininaya (myininaya):

first equation*

myininaya (myininaya):

\[\cot(x)=\sqrt{3} \text{ when } x=30^o \text{ or } x=210^o\]

OpenStudy (anonymous):

ahh gotcha on the function chart cot √3 = pi/6, 7pi/6

myininaya (myininaya):

\[\sin(x)=-\frac{\sqrt{2}}{2} \text{ when } x=225^o \text{ or } x=315^o\]

OpenStudy (anonymous):

thank you for clearing that up for me myininaya :) this damn book is worthless

myininaya (myininaya):

unit circle is your friend commit it to memory

OpenStudy (anonymous):

i even have the unit of circle chart, but its an online class so im teaching myself everything... or trying to :S

myininaya (myininaya):

do you see the pattern to remember the unit circle?

OpenStudy (anonymous):

dont so much remember the patter but i understand now what i was doing wrong :D

myininaya (myininaya):